How many spherical 4’ diameter balloons are required to lift a 175-lb person and 25 pounds of equipment into the air at standard sea level conditions? Assume that each balloon has a mass of 250 grams.
Barnard,
Answer #1

first

175 lb = 79.37 kg

25 pounds = 11.34

total weight to lift is = 79.37 + 11.34 = 90.71 kg

then the weight of balloon is = 0.25 kg

90.71 = 0.25 * n

n = 362.84